# AutoPresent: Designing Structured Visuals from Scratch

### v1

@@ 1 Introduction
To enable controlled and structured slide generation, we propose to create slides using program generation, where a model first generates a program from the natural language instruction, and then the program is executed to get the slide. We apply this approach to large language models (LLMs; LlaMa [11], GPT-4o [2]) and vision-language models (VLMs; LlaVa [49]). As illustrated in Figure 1, given a natural language instruction, the model first generates a Python program and then executes it to obtain a PPTX slide. We find that small models such as LlaMa (8B) and LlaVa (7B) are often unable to produce executable code. While GPT-4o can produce reasonable slides, it still exhibits a substantial gap in design quality compared to human-generated slides (§5). By further conducting iterative refinement, we find that models can self-refine and further improve slide quality. We also find that code generation approaches substantially outperform end-to-end image generation methods (Stable Diffusion [31], Dall-E [8]).
@@ 5.1 Experimental Setup
For code generation approaches, we sample $n=3$ responses and iteratively go through them, using the first successfully executing program as the final output of the model. If none of the $n$ responses execute successfully, we count it as an execution failure. In addition to AutoPresent4.2), we benchmark several LMs out-of-the-box, including open-weights Llama 3.1 (8B, Instruct), the vision-language LlaVa v1.5 model (with a Vicuna-7B-v1.5 LM backbone); and the proprietary GPT-4o model (the gpt-4o-2024-08-06 checkpoint).
@@
Code Generation w/o Library
@@
Code Generation w/ Expert-Designed Library
@@ 5.2 Quantitative Results and Analysis
Effect of LibrarySlidesLib brings observable gains in LlaMa and LlaVa in all three scenarios by at most $34.0$ points; and similarly increases the strong GPT-4o performance across scenarios, especially when no images are provided. This suggests the benefits of generating more modular and concise programs for structured visual design.
@@ 5.5 Result after Iterative Refinement
Method Detailed + Images Detailed Only High-Level
GPT-4o 58.0 56.3 58.5
Refinement 59.5 59.5 59.8
Table 6: Overall scores after applying refinement in the three scenarios, demonstrating that refinement boosts performance in all three scenarios, especially the detailed instructions only task.
Finally, as shown by Table 6, we find that refinement improves model performance on all three challenges, especially in the detailed instruction only setup by $3.2$ points. We present representative cases in Figure 5, which indicates that refinement can improve content layout and detailed controls on coloring and sizing.

### v2

@@ 1 Introduction
To enable controlled and structured slide generation, we propose to create slides using program generation, where a model first generates a program from the natural language instruction, and then the program is executed to get the slide. We apply this approach to large language models (LLMs; LlaMa [11], DeepseekCoder [16], CodeLlama [33], GPT-4o [2]) and vision-language models (VLMs; LlaVa [51]). As illustrated in Figure 1, given a natural language instruction, the model first generates a Python program and then executes it to obtain a PPTX slide. We find that small models such as LlaMa (8B) and LlaVa (7B) are often unable to produce executable code. While GPT-4o can produce reasonable slides, it still exhibits a substantial gap in design quality compared to human-generated slides (§5). By further conducting iterative refinement, we find that models can self-refine and further improve slide quality. We also find that code generation approaches substantially outperform end-to-end image generation methods (Stable Diffusion [32], Dall-E [8]).
@@ 5.1 Experimental Setup
For code generation approaches, we sample $n=3$ responses and iteratively go through them, using the first successfully executing program as the final output of the model. If none of the $n$ responses execute successfully, we count it as an execution failure. In addition to AutoPresent4.2), we benchmark several LMs out-of-the-box, including open-weights Llama 3.1 (8B, Instruct), the code generation models DeepseekCoder-7B-v1.5 and CodeLlaMa-7B-Instruct, the vision-language LlaVa v1.5 model (with a Vicuna-7B-v1.5 LM backbone); and the proprietary GPT-4o model (the gpt-4o-2024-08-06 checkpoint).
@@
Code Generation w/o SlidesLib
@@
Code Generation w/ SlidesLib
@@ 5.2 Quantitative Results and Analysis
Effect of SlidesLibSlidesLib brings observable gains in LlaMa and LlaVa in all three scenarios by at most $34.0$ points; and similarly increases the strong GPT-4o performance across scenarios, especially when no images are provided. This suggests the benefits of generating more modular and concise programs for structured visual design.
@@ 5.5 Result after Iterative Refinement
Iteration Detailed + Images Detailed Only High-Level
0 58.0 56.3 58.5
1 59.5 59.5 59.8
2 59.3 60.1 61.3
3 60.1 59.4 61.4
Table 6: Overall scores after applying multi-rounds of refinement in the three scenarios, demonstrating that refinement boosts performance in all three scenarios.
Finally, as shown by Table 6, we find that refinement improves model performance on all three challenges. By doing an ablation on the round of iterations, we find that while continued refinement often increases the scores, the first iteration usually gives the biggest performance improvement. We present representative cases after doing one round of refinement in Figure 5, which indicates that refinement can improve content layout and detailed controls on coloring and sizing.
@@ A.2 Slide Deck Source
A.3 Slide Deck Statistics Per Domain
The average images per domain and average text blocks per domain are shown in Figure 6.
Figure 6: SlidesBench statistics on different domains.
@@ B.2 SlidesLib Usage Example
Figure 10: The percentage of each action taken by different models.
B.3 SlidesLib Usage Percentage
The percentage of each action taken by GPT-4o, AutoPresent, and Llama-3.1 in all 3 scenarios are reported in Figure 10. On average, the most common actions are add_text (36.3%), add_image (20.3%), and add_title (13.5%).